# URLs Help, Topics Guide: Understanding Metadata Utility in LLM Training

### v1

@@ Abstract
Large Language Models (LLMs) are commonly pretrained on vast corpora of text without utilizing contextual metadata such as source, quality, or topic, leading to a context-free learning paradigm. While recent studies suggest that adding metadata like URL information as context (i.e., auxiliary inputs not used in the loss calculation) can improve training efficiency and downstream performance, they offer limited understanding of which types of metadata are truly effective and under what conditions. In this work, we conduct a systematic evaluation and find that not all metadata types contribute equally. Only URL context speeds up training, whereas quality scores and topic/format domain information offer no clear benefit. Furthermore, the improved downstream performances of URL conditioning emerge only when longer prompts are used at inference time. In addition, we demonstrate that context-aware pretraining enables more controllable generation than context-free pretraining, in a classifier-free guidance fashion. Although topic and format metadata do not accelerate training, they are effective for steering outputs, offering human-interpretable control over generation.
@@ 1 Introduction
Large Language Models (LLMs) have become increasingly integrated into a wide range of real-world applications, prompting both researchers and practitioners to explore optimal strategies for their development. A critical component of this process is the pretraining phase, which serves as the foundational stage in which models acquire general-purpose linguistic and world knowledge from vast corpora of text. Even with a simple next-token prediction loss, the scale of data and model size enables emergent generalization of knowledge [1].
@@
While prior studies [2, 3] suggest that incorporating URLs as context (auxiliary inputs not used in the loss calculation) during training can accelerate pretraining and enhance downstream performance, recent findings by Higuchi et al. [4] caution that this is not always the case. Existing work lacks a comprehensive understanding of which types of metadata are most impactful and under what conditions they offer the greatest benefit. Our study aims to fill this gap, providing empirical evidence based on real-world experiments.
Fundamentally, context-aware pretraining endowed LMs with the ability to perform conditional autoregressive generation. This conditional generation capability, developed through exposure to metadata, can potentially be leveraged to steer the model’s behavior using a classifier-free guidance approach [5], i.e., the distinction between context-conditioned and context-free generation can be further amplified to highlight the influence of contextual information. We aim to investigate whether context-aware pretraining can be utilized for more controllable generation than context-free pretraining.
@@ 2 Related Work
Allen-Zhu and Li [2] demonstrate through controlled synthetic experiments that adding a special token to high-credibility data can be beneficial. They further hypothesize that appending domain names (e.g., wikipedia.org) to each document in the pretraining corpus would yield similar gains. This hypothesis is experimentally validated by Gao et al. [3], who report a 33% speedup in pretraining simply by prepending URL domains. However, to ensure the model remains effective on standard, domain-free text, they introduce a "cool-down" phase using non-contextual data. Following this, Zhu et al. [6] formally proves that context-enhanced learning can be exponentially more sample-efficient when the model is capable of in-context learning, that is, it can effectively utilize contextual information presented during inference to perform tasks. More recently, Higuchi et al. [4] caution that metadata conditioning is not universally effective. Buidling up on the experimental setup of [2], they found that metadata conditioning only helps when the input prompt is sufficiently long to reveal latent semantics. With shorter prompts, it can even degrade performance. Most prior studies conduct experiments solely on synthetic datasets, where the metadata consists of the datasets’ production rules. However, in real-world scenarios, metadata cannot fully capture the complexity of natural language. Therefore, we aim to explore the extent to which such metadata can actually be helpful.
@@
In LLMs, next-token generation is essentially a sampling process from a probability distribution over the entire vocabulary. This makes it possible to steer the generation by modifying the token probability distribution. Liu et al. [7] guide the generation of a base LLM by leveraging the difference between the outputs of an expert model and an anti-expert model. Li et al. [8] adjust the sampling distribution by down-weighting tokens favored by an amateur model. Sanchez et al. [9] introduced the idea of classifier-free guidance (CFG)—originally used in text-to-image generation—to text-only generation. Building on this, He et al. [10] applied CFG to steer LLMs for personalization purposes, showing potential for controllable personalization. However, all of these approaches rely on context-free pretrained LLMs. In this work, we investigate can context-conditioned pretrained LLMs enable more effective context steering.
@@ 3.1 Context-Conditioned Pretraining
Context is introduced at the sequence level and is specific to each document. To ensure the model remains capable of handling context-free text at inference time, we uniformly interleave context-prepended and context-free documents during pretraining, using a 90%:10% ratio. The context enhanced corpus is illustrated in Figure 1, where one document can be split into several sequences and contexts are prepended to each of them. Unlike the MeCo approach [3], which applies 90% context-prepended training followed by a 10% context-free cooldown phase, our uniform mixture strategy allows any intermediate checkpoint to be immediately used for context-free inference.
@@ 3.2 Context-Aware Generation
    Context-Guided sampling: following the classifier-free guidance in diffusion models [5], we amplify the impact of the context with the following sampling probability:
@@ 4 Experiments
Model. We adopt the Llama model architecture [11] with 16 layers, a hidden size of 2048, a sequence length of 4096, and a batch size of 504 (resulting in 2.06 million tokens). The model has 1.5 billion parameters. We follow the Cosine learning schedule, applying 2000 warmup steps. AdamW optimizer is used with regularization strength 0.1 [12]. A max learning rate of 3e-4 is applied and cools down to learning rate 3e-5 at the end of training. To train the models, we use the Megatron-LM framework [13].
Dataset. We use FineWeb-Edu dataset [14], which is a high-quality English-only dataset. The dataset has a lot of meta data available, such as URL source, quality score and token counts per document. We sub-sample FineWeb-Edu and tokenize it using Nemo tokenizer [^3] https://mistral.ai/news/mistral-nemo, with two additionally introduced tokens. Throughout training, we randomly sample 100B tokens.
Evaluation Benchmarks. As standard practice, we evaluate models on general knowledge understanding using LM-Eval-Harness developed by Gao et al. [15]. The benchmarks used are Arc-Easy [16], Arc-Challenge [16], CommonSense QA (CSQA, 17), MMLU [18], PIQA [19], Social IQA (SIQA, 20), HellaSwag (HS, 21), Lambada (LBD, 22) and Winogrande (WG, 23).
@@
    • MeCo: A two-phase approach proposed by Gao et al. [3]—90% of training is URL-conditioned, followed by a 10% cooldown phase using standard (unconditioned) data.
@@
    • Domain Information (DI): Each document is labeled with topic and format domains derived from WebOrganizer [24], where topic and format domains are returned by pretrained classifiers. There are 24 categories per taxonomy. In total, we have 576 different Domain Information types.
@@ 4.1.1 Training Speed-up with URL Conditioning
Does the acceleration translate into downstream performance as well? With five-shot evaluation, we see better performance for the URL-conditioned model in every task. For the 9-task average, the URL-conditioned model can match the performance of the standard pretrained model on 100B tokens with only 60B tokens, achieving a 40% significant acceleration. The acceleration in terms of downstream performance is greater than training loss.
@@ 4.1.2 Downstream Performance Gains from Longer Prompts
When does URL conditioning work? To better understand the impact of context conditioning, we perform a systematic evaluation across downstream tasks by comparing zero-shot and five-shot performance. We notice a very interesting phenomenon: The URL-conditioned pretrained model shows increased downstream evaluation performance with five-shot evaluation, but not zero-shot. For the rest, we do not see substantial improvement over standard pretraining. Our results confirm recent findings from [4], where longer prompts are needed to infer latent semantics and thus to help evaluation.
Conflicting signals of different metadata types. Interestingly, the benefits of URL conditioning are negated when additional metadata is introduced as context. For example, combining quality scores with URL conditioning appears to introduce conflicting signals, leading to downstream performance that is even lower than that of the standard pretrained model. This suggests that the latent clusters inferred by the language model from the URLs may differ from the predefined clusters we assigned. Notably, we did not manage to reproduce the results of the MeCo baseline. Possible reasons could be different datasets and models, and that LM-Eval-Harness [15] is used as our evaluation framework is used in our experiments, while OLMES [25] was used in the MeCo paper. Nonetheless, we still observe the speeding-up effect of URL prepending.
@@ 4.2.2 Context-Guided Generation
[…]
@@ 5 Conclusion
Acknowledgment. This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a06 on Alps.

### v2

@@ Abstract
Large Language Models (LLMs) are commonly pretrained on vast corpora of text without utilizing contextual metadata such as source, quality, or topic, leading to a context-free learning paradigm. While recent studies suggest that adding metadata like URL information as context (i.e., auxiliary inputs not used in the loss calculation) can improve training efficiency and downstream performance, they offer limited understanding of which types of metadata are truly effective and under what conditions. In this work, we conduct a systematic evaluation and find that not all metadata types contribute equally. Only URL context speeds up training, whereas quality scores and topic/format domain information offer no clear benefit. Furthermore, the improved downstream performances of URL conditioning emerge only when longer prompts are used at inference time. In addition, we demonstrate that context-aware pretraining enables more controllable generation than context-free pretraining, in a classifier-free guidance fashion. Although topic and format metadata do not accelerate training, they are effective for steering outputs, offering human-interpretable control over generation. Our codes are available at this link.
@@ 1 Introduction
Large Language Models (LLMs) have become increasingly integrated into a wide range of real-world applications, prompting both researchers and practitioners to explore optimal strategies for their development. A critical component of this process is the pretraining phase, which serves as the foundational stage in which models acquire general-purpose linguistic and world knowledge from vast corpora of text. Even with a simple next-token prediction loss, the scale of data and model size enables emergent generalization of knowledge [du2024understanding].
@@
While prior studies [allenzhu2024physicslanguagemodels33, gao2025metadataconditioningaccelerateslanguage] suggest that incorporating URLs as context (auxiliary inputs not used in the loss calculation) during training can accelerate pretraining and enhance downstream performance, recent findings by higuchi2025doesmetadataconditioningnot caution that this is not always the case. Existing work lacks a comprehensive understanding of which types of metadata are most impactful and under what conditions they offer the greatest benefit. Our study aims to fill this gap, providing empirical evidence based on real-world experiments.
Fundamentally, context-aware pretraining endowed LMs with the ability to perform conditional autoregressive generation. This conditional generation capability, developed through exposure to metadata, can potentially be leveraged to steer the model’s behavior using a classifier-free guidance approach [ho2022classifierfreediffusionguidance], i.e., the distinction between context-conditioned and context-free generation can be further amplified to highlight the influence of contextual information. We aim to investigate whether context-aware pretraining can be utilized for more controllable generation than context-free pretraining.
@@ 2 Related Work
allenzhu2024physicslanguagemodels33 demonstrate through controlled synthetic experiments that adding a special token to high-credibility data can be beneficial. They further hypothesize that appending domain names (e.g., wikipedia.org) to each document in the pretraining corpus would yield similar gains. This hypothesis is experimentally validated by gao2025metadataconditioningaccelerateslanguage, who report a 33% speedup in pretraining simply by prepending URL domains. However, to ensure the model remains effective on standard, domain-free text, they introduce a "cool-down" phase using non-contextual data. Following this, zhu2025powercontextenhancedlearningllms formally proves that context-enhanced learning can be exponentially more sample-efficient when the model is capable of in-context learning, that is, it can effectively utilize contextual information presented during inference to perform tasks. More recently, higuchi2025doesmetadataconditioningnot caution that metadata conditioning is not universally effective. Buidling up on the experimental setup of [allenzhu2024physicslanguagemodels33], they found that metadata conditioning only helps when the input prompt is sufficiently long to reveal latent semantics. With shorter prompts, it can even degrade performance. Most prior studies conduct experiments solely on synthetic datasets, where the metadata consists of the datasets’ production rules. However, in real-world scenarios, metadata cannot fully capture the complexity of natural language. Therefore, we aim to explore the extent to which such metadata can actually be helpful.
@@
In LLMs, next-token generation is essentially a sampling process from a probability distribution over the entire vocabulary. This makes it possible to steer the generation by modifying the token probability distribution. liu-etal-2021-dexperts guide the generation of a base LLM by leveraging the difference between the outputs of an expert model and an anti-expert model. li-etal-2023-contrastive adjust the sampling distribution by down-weighting tokens favored by an amateur model. sanchez2023staytopicclassifierfreeguidance introduced the idea of classifier-free guidance (CFG)—originally used in text-to-image generation—to text-only generation. Building on this, he2025contextsteeringcontrollablepersonalization applied CFG to steer LLMs for personalization purposes, showing potential for controllable personalization. However, all of these approaches rely on context-free pretrained LLMs. In this work, we investigate can context-conditioned pretrained LLMs enable more effective context steering.
@@ 3.1 Context-Conditioned Pretraining
Context is introduced at the sequence level and is specific to each document. To ensure the model remains capable of handling context-free text at inference time, we uniformly interleave context-prepended and context-free documents during pretraining, using a 90%:10% ratio. The context enhanced corpus is illustrated in Figure 1, where one document can be split into several sequences and contexts are prepended to each of them. Unlike the MeCo approach [gao2025metadataconditioningaccelerateslanguage], which applies 90% context-prepended training followed by a 10% context-free cooldown phase, our uniform mixture strategy allows any intermediate checkpoint to be immediately used for context-free inference.
@@ 3.2 Context-Aware Generation
    Context-Guided sampling: following the classifier-free guidance in diffusion models [ho2022classifierfreediffusionguidance], we amplify the impact of the context with the following sampling probability:
@@ 4 Experiments
Model. We adopt the Llama model architecture [grattafiori2024llama3herdmodels] with 16 layers, a hidden size of 2048, a sequence length of 4096, and a batch size of 504 (resulting in 2.06 million tokens). The model has 1.5 billion parameters. We follow the Cosine learning schedule, applying 2000 warmup steps. AdamW optimizer is used with regularization strength 0.1 [loshchilov2019decoupledweightdecayregularization]. A max learning rate of 3e-4 is applied and cools down to learning rate 3e-5 at the end of training. To train the models, we use the Megatron-LM framework [narayanan2021efficientlargescalelanguagemodel].
Dataset. We use FineWeb-Edu dataset [lozhkov2024fineweb-edu], which is a high-quality English-only dataset. The dataset has a lot of meta data available, such as URL source, quality score and token counts per document. We sub-sample FineWeb-Edu and tokenize it using Nemo tokenizer [^3] https://mistral.ai/news/mistral-nemo, with two additionally introduced tokens. Throughout training, we randomly sample 100B tokens.
Evaluation Benchmarks. As standard practice, we evaluate models on general knowledge understanding using LM-Eval-Harness developed by eval-harness. The benchmarks used are Arc-Easy [clark2018think], Arc-Challenge [clark2018think], CommonSense QA (CSQA, talmor2018commonsenseqa), MMLU [hendrycks2020measuring], PIQA [bisk2020piqa], Social IQA (SIQA, sap2019socialiqa), HellaSwag (HS, zellers2019hellaswag), Lambada (LBD, paperno2016lambada) and Winogrande (WG, sakaguchi2021winogrande).
@@
    • MeCo: A two-phase approach proposed by gao2025metadataconditioningaccelerateslanguage—90% of training is URL-conditioned, followed by a 10% cooldown phase using standard (unconditioned) data.
@@
    • Domain Information (DI): Each document is labeled with topic and format domains derived from WebOrganizer [wettig2025organizewebconstructingdomains], where topic and format domains are returned by pretrained classifiers. There are 24 categories per taxonomy. In total, we have 576 different Domain Information types.
@@ 4.1.1 Training Speed-up with URL Conditioning
Does the acceleration translate into downstream performance as well? With five-shot evaluation, we see better performance for the URL-conditioned model in every task. For the 9-task average, the URL-conditioned model can match the performance of the standard pretrained model on 100B tokens with only 60B tokens, achieving a 40% significant acceleration, as shown in Figure 4. The acceleration in terms of downstream performance is greater than training loss.
@@
Why only URL helps? We have already controlled for the confounding factor — the length of metadata. However, it remains unclear why only the URL contributes meaningfully. To gain insight into this, we visualized the attention patterns of the document texts toward different parts of their corresponding contexts in Figure 5. Except for the URL-conditioned model, the other models fail to allocate substantial attention to the informative parts of the contexts in the early layers. We hypothesize that the early-layer attention to meaningful URL components may play a key role in the effectiveness of URL prepending, even if later layers focus on less informative tokens. Moreover, it is important to recognize that URLs inherently encode both domain information and document quality, which may differ from human-assigned DI and QS labels. This distinction likely also contributes to the superior performance of the URL feature.
@@
Figure 5: Average attention to different parts of different prepended contexts. Take https://en.wikipedia.org/wiki/Metadata#Standards as an example. en.wikipedia.org is the URL domain and /wiki/Metadata#Standards is the URL suffix. More details in Appendix C.1.2.
@@ 4.1.2 Downstream Performance Gains from Longer Prompts
When does URL conditioning work? To better understand the impact of context conditioning, we perform a systematic evaluation across downstream tasks by comparing zero-shot and five-shot performance. We notice a very interesting phenomenon: The URL-conditioned pretrained model shows increased downstream evaluation performance with five-shot evaluation, but not zero-shot. For the rest, we do not see substantial improvement over standard pretraining. Our results confirm recent findings from higuchi2025doesmetadataconditioningnot, where longer prompts are needed to infer latent semantics and thus to help evaluation.
Conflicting signals of different metadata types. Interestingly, the benefits of URL conditioning are negated when additional metadata is introduced as context. For example, combining quality scores with URL conditioning appears to introduce conflicting signals, leading to downstream performance that is even lower than that of the standard pretrained model. This suggests that the latent clusters inferred by the language model from the URLs may differ from the predefined clusters we assigned. Notably, we did not manage to reproduce the results of the MeCo baseline. Possible reasons could be different datasets and models, and that LM-Eval-Harness [eval-harness] is used as our evaluation framework is used in our experiments, while OLMES [gu-etal-2025-olmes] was used in the MeCo paper. Nonetheless, we still observe the speeding-up effect of URL prepending.
@@ 4.2.2 Context-Guided Generation
https://reddit.com/ 4 things, its cuisine, its fashion, its nightlife and its cinema. And I don’t mean that only. [Uncaptioned image] What if I told you that there is a community of actors who actually act out the scenes of French movies? This community is called REDdit!
@@ 5 Conclusion
Acknowledgment. This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a06 on Alps. DF would like to thank Diba Hashemi for helpful discussions.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our main text corresponds to what we claim in the abstract and introduction, i.e., what benefits does context-aware pretraining bring in pretraining and inference time, respectively.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See our discussions in the conclusion section as well as in Appendix A.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We detailed our tokenization process and metadata used. Moreover, the training details are also specified in Section 4.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We will make our training codes public when the paper is made public. The dataset we used is public.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specified our dataset, hyperparameter choices, and training details in Section 4.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We reported 3-seed run results with both mean and standard deviation in Section 4.2.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We specified in the Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We adhere to the Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See Appendix B.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We are not releasing new models, but rather providing insights to assist model pretraining.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cited the dataset, model training framework, codebase and tokenizer.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We clearly described our usage of LLM as a judge, to grade continuations from different models.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.